An automobile engine can run on a mixture of gasoline and a substitute fuel. If gas costs $3.45 per gallon and the substitute fuel costs $2 per gallon, what percent of a mixture must be substitute fuel to bring the cost down to $2.58 per gallon?

Answer by ankor@dixie-net.com(22740)

If gas costs $3.45 per gallon and the substitute fuel costs $2 per gallon, what percent of a mixture must be substitute fuel to bring the cost down to $2.58 per gallon?
:
We can make 10 gal of the mixture and find the percent from that
Let x = amt of 3.45 fuel
then
(10-x) = amt of 2.00 fuel
:
3.45x + 2(10-x) = 2.58(10)
3.45x + 20 - 2x = 25.8
3.45x - 2x = 25.8 - 20
1.45x = 5.8
x = 5.8/1.45
x = 4 gal of 3.45 fuel
then
10-4 = 6 gal of $2 fuel
Obviously that is 60%